### v1

@@ 3.3 Coming-down-from-infinity property
In this subsection, we establish the coming-down-from-infinity property for the coalescent process $X$ under the condition (H2’) introduced below.
3.3.1 Condition (H2’)
We now introduce the condition required to establish the coming-down-from-infinity property. Define for $n\geq 1$, $0\leq s\leq h(q_{n})$ and $1\leq m\leq q_{n}(s)$,
@@ 3.3.1 Condition (H2’)
By Theorem 3.14, this quantity also equals ${\mathbb{P}}_{n}(X_{s-1}\leq m-1\,|\,X_{s}=m)$ for any coalescent process $X$ based at a height strictly greater than $s$.
The condition is formulated as follows:
  1. (H2’)
    For every $\delta>0$, there exists $\varepsilon>0$ such that uniformly for all $n\geq 1$ and $\delta n\leq s\leq h(q_{n})-\delta n$,
    $p_{n}\!\left(\lfloor(\log n)^{1+\varepsilon}\rfloor,s\right)\gtrsim\frac{(\log n)^{2(1+\varepsilon)}}{n}.$ (35)
Remark 3.22.
Condition (H2’) provides a lower bound of the form
$p_{n}(m,s)\gtrsim\frac{m}{n^{2}}$ (36)
when $m=\lfloor(\log n)^{1+\varepsilon}\rfloor$. As will be shown in Lemma 3.23, in order for (H2’) to hold, it suffices that for each $n\geq 1$, there exists some $m^{\prime}\geq(\log n)^{1+\varepsilon}$ for which the bound (36) holds; the bound then automatically propagates to all smaller $m\leq m^{\prime}$.
The following lemma formalizes this monotonicity property and will be used to control the coalescence rate in the proof of the coming-down-from-infinity property.
@@
For any $n\geq 1$, $1\leq s\leq h(q_{n})$, and $1\leq m^{\prime}\leq m\leq q_{n}(s)$,
$p_{n}(m^{\prime},s)\;\geq\;\frac{(m^{\prime})_{2}}{(m)_{2}}\,p_{n}(m,s).$
Proof 3.24.
Recall the notion of coalescent event from (13). Using this, we have the disjoint decomposition:
$\big\{\#\{\mathfrak{p}(u^{s}_{1}),\ldots,\mathfrak{p}(u^{s}_{m})\}\leq m-1\big\}=\bigcup_{\begin{subarray}{c}f\in M^{s}:\text{Dom}(f)=\{1,\ldots,m\}\\ \text{Range}(f)\leq m-1\end{subarray}}\text{Coal}^{s}(f).$
On any event $\mathrm{Coal}^{s}(f)$ appearing above, there is at least one collision of parents: there exist indices $i<j\leq m$ such that $\mathfrak{p}(u^{s}_{i})=\mathfrak{p}(u^{s}_{j})$. Note that conditionally on $\mathrm{Coal}^{s}(f)$, the $m^{\prime}$ individuals $\{u^{s}_{1},\ldots,u^{s}_{m^{\prime}}\}$ are obtained by choosing an $m^{\prime}$-subset uniformly from $\{u^{s}_{1},\ldots,u^{s}_{m}\}$. Thus the conditional probability that the colliding pair $\{i,j\}$ is entirely contained in $\{1,\ldots,m^{\prime}\}$ is $\frac{(m^{\prime})_{2}}{(m)_{2}}$. Consequently,
$\mathbb{P}_{n}\big(\#\{\mathfrak{p}(u^{s}_{1}),\ldots,\mathfrak{p}(u^{s}_{m})\}\leq m-1\,|\,\text{Coal}^{s}(f)\big)\geq\frac{(m^{\prime})_{2}}{(m)_{2}}.$
This leads to the conclusion.
Corollary 3.25.
@@
Assume (H2’). For every $\delta>0$, there exists $\varepsilon>0$ such that uniformly for all $n\geq 1$, $\delta n\leq s\leq h(q_{n})-\delta n$ and $2\leq m\leq(\log n)^{1+\varepsilon}$,
$p_{n}\!\left(m,s\right)\gtrsim\frac{m^{2}}{n}.$ (37)
3.3.2 Coming-down-from-infinity property
Fix two sequences $h^{*}(n)$ and $h_{*}(n)$, $n\geq 1$, such that
$0<h_{*}(n)<h^{*}(n)<h(q_{n}),\qquad n\geq 1,$
and assume there exist constants $0<H_{*}<H^{*}<h(\ell)$ such that
$\frac{h_{*}(n)}{n}\to H_{*}\quad\text{and}\quad\frac{h^{*}(n)}{n}\to H^{*}.$ (38)
In particular, (38) implies that there exists $\delta>0$ such that, for all $n\geq 1$,
$\delta n\leq h_{*}(n)<h^{*}(n)\leq h(q_{n})-\delta n\quad\text{and}\quad h^{*}(n)-h_{*}(n)\geq\delta n.$ (39)
We consider the coalescent process $X_{j}=X_{j}(n,h^{*}(n))$ based at height $h^{*}(n)$ and focus on its value at height $h_{*}(n)$:
$N_{n}:=X_{h_{*}(n)}(n,h^{*}(n)).$
For brevity, we write $h^{*}=h^{*}(n)$ and $h_{*}=h_{*}(n)$ when $n$ is clear from the context.
Theorem 3.26 (Coming-down-from-infinity property).
Assume (H1) and (H2’). Then the sequence $(r_{n})$ is tight, i.e.,
$\lim_{M\to\infty}\,\limsup_{n\to\infty}\,{\mathbb{P}}_{n}(N_{n}>M)=0.$
Remark 3.27.
This can be regarded as a sequential version of the coming-down-from-infinity property, since the bound $M$ must be chosen independently of $n$.
Proof 3.28.
For each $M\geq 1$, define the hitting time
$T_{M}:=T_{M}(n):=\sup\{0\leq j\leq h^{*}(n):X_{j}\leq M\}.$
Clearly,
$\{N_{n}>M\}=\{T_{M}<h_{*}\}\subset\{T_{M}\leq h^{*}-\delta n\}.$
Thus, it suffices to prove
$\lim_{M\to\infty}\limsup_{n\to\infty}{\mathbb{P}}_{n}(T_{M}\leq h^{*}-\delta n)=0,$ (40)
by (39). Hence it is enough to show that for any $\eta>0$, there exists $M\geq 1$ such that, for all sufficiently large $n$,
${\mathbb{P}}_{n}\Big(T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\leq h^{*}-\tfrac{2}{3}\delta n\Big)<\tfrac{\eta}{2},$ (41)
${\mathbb{P}}_{n}\Big(T_{M}\leq T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-\tfrac{1}{3}\delta n\,\Big|\,T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\tfrac{2}{3}\delta n\Big)<\tfrac{\eta}{2}.$ (42)
Proof 3.29 (Proof of (42)).
Define
@@ 3.3.2 Coming-down-from-infinity property
$T_{m_{2}}-T_{m_{1}}=\sum_{m=m_{1}+1}^{m_{2}}\zeta_{m}.$ (43)
By the strong Markov property, conditionally on $T_{m}=j$ and $X_{T_{m}}=m$,
$\zeta_{m}\overset{d}{=}\inf\{1\leq i\leq j:Y_{i}=1\},$ (44)
@@
By (37), there exist $C^{*}>0$ such that for $n\geq 1$, $h_{*}\leq i\leq h^{*}$ and $2\leq m\leq\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor$,
${\mathbb{P}}(Y_{i}=1)\geq C^{*}\frac{m^{2}}{n}.$ (45)
@@
For any $2\leq m<\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor$, conditionally on $T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\frac{2}{3}\delta n$, we have the stochastic bound
$\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)^{+}\;\leq_{\rm sto}\;{\rm Geometric}(C^{*}m^{2}/n).$ (46)
@@
If $T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-T_{M}=\sum_{m=M+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{m}\geq\tfrac{1}{3}\delta n$, then
$\sum_{m=M+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)^{+}\geq\tfrac{1}{3}\delta n.$ (47)
Based on these lemmas, let $\gamma_{m}$ be independent ${\rm Geometric}(C^{*}m^{2}/n)$ variables. Then
@@
Since
${\mathbb{P}}\Big(\sum_{m=M+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\gamma_{m}\geq\tfrac{1}{3}\delta n\Big)\;\leq\;\frac{3}{\delta n}{\mathbb{E}}\Big[\sum_{m=M+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\gamma_{m}\Big]\;\leq\;\frac{c}{M},$
for some constant $c>0$, substituting into (48) gives (42) once $M$ is chosen sufficiently large.
Proof 3.32 (Proof of Lemma 3.30).
It suffices to show that conditionally on $\big(X_{j}:T_{m}\leq j\leq h^{*}\big)$ with $T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\frac{2}{3}\delta n$, $X_{T_{m}}=m$ and $\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}>0$,
$\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)\;\leq_{\rm sto}\;{\rm Geometric}(C^{*}m^{2}/n).$ (49)
@@
$T_{m}=T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-\sum_{i=j+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}$ $>h^{*}-\delta n+\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)$
@@
Then (49) follows from (44) and (45).
Proof 3.33 (Proof of Lemma 3.31).
Assume $\sum_{m=M+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{m}\geq\tfrac{1}{3}\delta n.$ Define
$m^{*}:=\sup\Big\{\,m\geq M+1:\ \sum_{j=m}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\geq\tfrac{1}{3}\delta n\,\Big\}.$
@@
$\sum_{j=m^{*}}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\geq\tfrac{1}{3}\delta n\quad\text{and}\quad\sum_{j=m}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}<\tfrac{1}{3}\delta n\ \text{for all }m>m^{*}.$
Therefore
$\sum_{m=m^{*}}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)^{+}$ $=\sum_{m=m^{*}}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{m}+\left(\frac{1}{3}\delta n-\sum_{j=m^{*}+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\zeta_{j}\right)^{+}$
$\geq\frac{1}{3}\delta n.$
This proves (47).
We now turn to the proof of (41). The main technical estimate we require is the following proposition.
@@
Assume (H1) and (H2’). Then there exist constants $c>0$ such that for any $n\geq 1$, $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}\leq s\leq h^{*}$ and $1\leq k\leq q_{n}(j)$ and , we have
${\mathbb{P}}_{n}\left(X_{s-{\frac{n}{(\log n)^{1+\frac{\varepsilon}{4}}}}}\leq\dfrac{k}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}}\,\Big|\,X_{s}=k\right)\geq 1-e^{-c(\log n)^{\frac{\varepsilon}{4}}}.$ (50)
Proof 3.35 (Proof of (41) assuming Proposition 3.34).
@@
$l_{i}=h^{*}-\dfrac{n\cdot i}{(\log n)^{1+\frac{\varepsilon}{4}}}$
for $i=0,\cdots,\lfloor\frac{2}{3}\delta(\log n)^{1+\frac{\varepsilon}{4}}\rfloor$. We use the following elementary fact:
@@
$X_{l_{i+1}}\leq\dfrac{X_{l_{i}}}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}},$
for all $i=0,\ldots,\lfloor\frac{2}{3}\delta(\log n)^{1+\frac{\varepsilon}{4}}\rfloor$, then $X_{\lfloor\frac{2}{3}\delta n\rfloor}<2(\log n)^{1+\frac{\varepsilon}{2}}$.

### v2

@@ 3.3 Coming-down-from-infinity property
In this subsection, we establish the CDFI property for the coalescent process $X$.
Fix two sequences $h^{*}(n)$ and $h_{*}(n)$, $n\geq 1$, such that
$0<h_{*}(n)<h^{*}(n)<h(q_{n}),$ $\forall\,n\geq 1,$ (3.27)
$\frac{h_{*}(n)}{n}\to H_{*}\quad\text{and}\quad\frac{h^{*}(n)}{n}\to H^{*},$ $n\to\infty,$
with some constants $0<H_{*}<H^{*}<h(\ell)$. In particular, (3.27) together with (H1) implies that there exists $\delta>0$ such that, for all $n\geq 1$,
$\delta n\leq h_{*}(n)<h^{*}(n)\leq h(q_{n})-\delta n\quad\text{and}\quad h^{*}(n)-h_{*}(n)\geq\delta n.$ (3.28)
We consider the coalescent process $X_{j}=X_{j}(n,h^{*}(n))$ based at height $h^{*}(n)$ and focus on its value at height $h_{*}(n)$, namely $X_{h_{*}(n)}$. For brevity, we write $h^{*}=h^{*}(n)$ and $h_{*}=h_{*}(n)$ when $n$ is clear from the context.
Theorem 3.13 (CDFI property).
Assume (H1) and (H2a). Then the sequence $(X_{h_{*}(n)})$ is tight, i.e.
$\lim_{M\to\infty}\,\limsup_{n\to\infty}\,{\mathbb{P}}_{n}(X_{h_{*}}>M)=0.$
Remark.
This can be regarded as a sequential version of the CDFI property, since the bound $M$ must be chosen independently of $n$.
Proof.
For each $M\geq 1$, define the hitting time
$T_{M}=T_{M}(n):=\sup\{0\leq j\leq h^{*}(n):X_{j}\leq M\}.$
Clearly,
$\{X_{h_{*}}>M\}=\{T_{M}<h_{*}\}\subset\{T_{M}\leq h^{*}-\delta n\}$
by (3.28). Hence it suffices to prove
$\lim_{M\to\infty}\limsup_{n\to\infty}{\mathbb{P}}_{n}(T_{M}\leq h^{*}-\delta n)=0.$ (3.29)
It is enough to show that for any $\eta>0$, there exists $M=M(\eta)\geq 1$ such that, for all sufficiently large $n$,
${\mathbb{P}}_{n}\Big(T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\leq h^{*}-\tfrac{1}{3}\delta n\Big)<\tfrac{\eta}{3},$ (3.30)
${\mathbb{P}}_{n}(T_{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\leq T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-\tfrac{1}{3}\delta n\,\Big|\,T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\tfrac{1}{3}\delta n\Big)<\tfrac{\eta}{3},$ (3.31)
${\mathbb{P}}_{n}\Big(T_{M}\leq T_{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}-\tfrac{1}{3}\delta n\,\Big|\,T_{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}>T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-\tfrac{1}{3}\delta n,\,T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\tfrac{1}{3}\delta n\Big)<\tfrac{\eta}{3}.$ (3.32)
3.3.1 Proof of (3.31) and (3.32)
In this part, we prove (3.31) and (3.32). The key step is to obtain a lower bound on the one-step transition probabilities of the process $X_{j}$ using (H2a). This allows us to stochastically dominate $T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-T_{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}$ and $T_{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}-T_{M}$ by sums of independent geometric random variables. Consequently, with high probability, both increments can be made arbitrarily small.
Define for $n\geq 1$, $0\leq s\leq h(q_{n})$ and $1\leq m\leq q_{n}(s)$,
@@
By Theorem 3.9, this quantity also equals ${\mathbb{P}}_{n}(X_{s-1}\leq m-1\,|\,X_{s}=m)$ for any coalescent process $X$ based at a height greater than $s$.
@@
Assume (H1) and (H2a). For every $\delta>0$, there exists $\varepsilon^{\prime}>0$ such that uniformly for all $n\geq 1$, $\delta n\leq s\leq h(q_{n})-\delta n$ and $2\leq m\leq(\log n)^{\frac{1+\varepsilon^{\prime}}{2}}$,
$p_{n}\!\left(m,s\right)\gtrsim\frac{m^{2}}{n}.$ (3.33)
Proof.
By (H2a), for each $\delta>0$ there exist $N=N(\delta)$ such that
$\tilde{\sigma}^{2}:=\inf_{n\geq N,\,\delta n\leq s\leq h(q_{n})-\delta n}{\mathbb{E}}_{n}\bigg[(\nu^{s}_{1})_{2}\mathbf{1}\bigg\{\nu_{1}^{s}\leq\frac{n}{(\log n)^{1+\varepsilon}}\bigg\}\bigg]>0.$ (3.34)
Define $K_{n}=\lfloor\frac{n}{(\log n)^{1+\varepsilon}}\rfloor$ and truncated coalescent events
$B_{i_{1}i_{2}}:=\big\{\mathfrak{p}(u^{s}_{i_{1}})=\mathfrak{p}(u^{s}_{i_{2}}),\,\xi(\mathfrak{p}(u^{s}_{i_{1}}))\leq K_{n}\big\},\qquad 1\leq i_{1}<i_{2}\leq m.$
Then
$\Big\{\#\{\mathfrak{p}(u^{s}_{i}):1\leq i\leq m\}\leq m-1\Big\}=\bigcup_{1\leq i_{1}<i_{2}\leq m}\big\{\mathfrak{p}(u^{s}_{i_{1}})=\mathfrak{p}(u^{s}_{i_{2}})\big\}\supset\bigcup_{1\leq i_{1}<i_{2}\leq m}B_{i_{1}i_{2}}.$ (3.35)
Applying the inclusion–exclusion principle, we get
By Corollary 3.3, for all $n\geq N$,
${\mathbb{P}}_{n}(B_{12})$ $=\frac{q_{n}(s-1)}{(q_{n}(s))_{2}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{2}\mathbf{1}\{\nu^{s-1}_{1}\leq K_{n}\}\big]\;\geq\;\frac{q_{n}(s-1)}{(q_{n}(s))_{2}}\,\tilde{\sigma}^{2},$
${\mathbb{P}}_{n}(B_{12}\cap B_{13})$ $=\frac{q_{n}(s-1)}{(q_{n}(s))_{3}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{3}\mathbf{1}\{\nu^{s-1}_{1}\leq K_{n}\}\big]\;\leq\;\frac{q_{n}(s-1)}{(q_{n}(s))_{3}}\,K_{n}{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{2}\big],$
${\mathbb{P}}_{n}(B_{12}\cap B_{34})$ $\leq\frac{q_{n}(s-1)}{(q_{n}(s))_{4}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{4}\mathbf{1}\{\nu^{s-1}_{1}\leq K_{n}\}\big]+\frac{(q_{n}(s-1))_{2}}{(q_{n}(s))_{4}}\,{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{2}(\nu^{s-1}_{2})_{2}\mathbf{1}\{\nu^{s-1}_{1},\nu^{s-1}_{2}\leq K_{n}\}\big]$
$\leq\frac{q_{n}(s-1)}{(q_{n}(s))_{4}}K_{n}^{2}{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{2}\big]+\frac{(q_{n}(s-1))_{2}}{(q_{n}(s))_{4}}\,K_{n}{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{2}\,\nu^{s-1}_{2}\big].$
We recall that ${\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{2}\,\nu^{s-1}_{2}\big]\leq\frac{q_{n}(s)}{q_{n}(s-1)-1}{\mathbb{E}}_{n}\!\big[(\nu^{s-1}_{1})_{2}\big]$ by (3.13). Hence by (H1), there exist constants $c_{i}=c_{i}(\delta,\tilde{\sigma},N)>0$ $(i=1,2,3)$ such that, for all $n\geq N$, $\delta n\leq s\leq h(q_{n})-\delta n$ and $1\leq m\leq q_{n}(s)$,
${\mathbb{P}}_{n}\Big(\bigcup_{1\leq i<j\leq m}B_{ij}\Big)\;\geq\;c_{1}\frac{m^{2}}{n}-c_{2}\frac{m^{3}}{n(\log n)^{1+\varepsilon}}-c_{3}\frac{m^{4}}{n(\log n)^{1+\varepsilon}}.$
Combining this with (3.35) and taking $\varepsilon^{\prime}\in(0,\varepsilon)$, we obtain (3.33). ∎
We are now ready to prove (3.31) and (3.32). Define
@@
$T_{m_{2}}-T_{m_{1}}=\sum_{m=m_{1}+1}^{m_{2}}\zeta_{m}.$ (3.36)
By the Markov property, conditionally on $\mathcal{G}_{j}$ with $T_{m}=j$ and $X_{T_{m}}=m$,
$\zeta_{m}\overset{d}{=}\inf\{1\leq i\leq j:Y_{i}=1\},$ (3.37)
@@
By (3.33) with $\varepsilon^{\prime}=2\varepsilon/3$, there exists constant $C^{*}>0$ such that for all $n\geq 1$, $h_{*}\leq i\leq h^{*}$ and $2\leq m\leq\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor$,
${\mathbb{P}}(Y_{i}=1)\geq C^{*}\,\frac{m^{2}}{n}.$ (3.38)
@@
Assume (H1) and (H2a). Conditionally on $T_{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}>T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-\tfrac{1}{3}\delta n$ and $T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\tfrac{1}{3}\delta n$, we have the stochastic bound:
$\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{j}\right)^{+}\;\leq_{\rm sto}\;{\rm Geometric}\bigg(\frac{C^{*}m^{2}}{n}\bigg),\quad\forall\,2\leq m<\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor.$ (3.39)
@@
If $T_{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}-T_{M}=\sum_{m=M+1}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{m}\geq\tfrac{1}{3}\delta n$, then
$\sum_{m=M+1}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{j}\right)^{+}\geq\frac{1}{3}\delta n.$ (3.40)
Proof of (3.32) assuming Lemmas (3.15) and (3.16).
Set
$A=A_{n}:=\left\{T_{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}>T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-\tfrac{1}{3}\delta n,\,T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\tfrac{1}{3}\delta n\right\}.$
Then
@@
for some constant $c>0$. This gives (3.32) once $M$ is chosen sufficiently ∎
Proof of (3.31).
Observe that ${\mathbb{P}}_{n}(X_{j}\leq m-1\mid X_{j-1}=m)={\mathbb{P}}_{n}\big(\#\{\mathfrak{p}(u^{j-1}_{i}):1\leq i\leq m\}\leq m-1\big)$ is increasing in $m$. In particular, for any $n\geq 1$, $\delta n\leq s\leq h(q_{n})-\delta n$ and $(\log n)^{\frac{3+2\varepsilon}{6}}\leq m<2(\log n)^{1+\frac{\varepsilon}{2}}$,
$p_{n}(m,s)\;\geq\;C^{*}\dfrac{(\log n)^{1+\frac{2}{3}{\varepsilon}}}{n}.$
Repeating the same argument as above yields
$\begin{split}&{\mathbb{P}}_{n}(T_{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\leq T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-\tfrac{1}{3}\delta n\,\Big|\,T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\tfrac{1}{3}\delta n\Big)\\ &\;\leq\;\frac{3}{\delta n}\sum_{m=\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor+1}^{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}\dfrac{n}{C^{*}(\log n)^{1+\frac{2}{3}{\varepsilon}}}\leq\frac{6}{C^{*}\delta(\log n)^{\frac{\varepsilon}{6}}},\end{split}$ (3.42)
which implies (3.31). ∎
Proof of Lemma 3.15.
It is enough to show that conditionally on $\big(X_{j}:T_{m}\leq j\leq h^{*}\big)$ with $X_{T_{m}}=m$, $T_{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}>T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}-\tfrac{1}{3}\delta n,\ T_{\lfloor 2(\log n)^{1+\frac{\varepsilon}{2}}\rfloor}>h^{*}-\tfrac{1}{3}\delta n$ and $\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{j}>0$,
$\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{j}\right)\;\leq_{\rm sto}\;{\rm Geometric}\bigg(\frac{C^{*}m^{2}}{n}\bigg).$ (3.43)
@@
Therefore the LHS of (3.43) is greater than $h_{*}$, and thus (3.43) follows from (3.37) and (3.38). ∎
Proof of Lemma 3.16.
Assume $\sum_{m=M+1}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{m}\geq\tfrac{1}{3}\delta n.$ Define
$m^{*}:=\sup\Big\{\,m\geq M+1:\ \sum_{j=m}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{j}\geq\tfrac{1}{3}\delta n\,\Big\}.$
@@
$\sum_{j=m^{*}}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{j}\geq\frac{1}{3}\delta n\quad\text{and}\quad\sum_{j=m}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{j}<\frac{1}{3}\delta n\quad\text{for all }m>m^{*}.$
Therefore, by separating the contributions of the cases $m=m^{*}$ and $m\geq m^{*}+1$, we derive that
$\sum_{m=m^{*}}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{m}\wedge\left(\frac{1}{3}\delta n-\sum_{j=m+1}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{j}\right)^{+}=\left(\frac{1}{3}\delta n-\sum_{j=m^{*}+1}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{j}\right)^{+}+\sum_{m=m^{*}+1}^{\lfloor(\log n)^{\frac{3+2\varepsilon}{6}}\rfloor}\zeta_{m}\geq\frac{1}{3}\delta n.$
This proves (3.40). ∎
3.3.2 Proof of (3.30)
We close this section with the proof of (3.30). The main technical estimate we require is the following proposition.
@@
Assume (H1) and (H2a). Then there exists constant $c>0$ such that for any $n$ sufficiently large, $h_{*}+\frac{n}{(\log n)^{1+\frac{\varepsilon}{8}}}\leq s\leq h^{*}$ and $1\leq k\leq q_{n}(s)$, we have
${\mathbb{P}}_{n}\left(X_{s-{\frac{n}{(\log n)^{1+\frac{\varepsilon}{8}}}}}\leq\dfrac{k}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}}\,\Big|\,X_{s}=k\right)\geq 1-c(\log n)^{-\frac{13}{24}\varepsilon}.$ (3.44)
Proof of (3.30) assuming Proposition 3.17.
@@
$l_{i}=h^{*}-\dfrac{n\cdot i}{(\log n)^{1+\frac{\varepsilon}{8}}}$
for $i=0,\cdots,\lfloor\frac{1}{3}\delta(\log n)^{1+\frac{\varepsilon}{8}}\rfloor$. We use the following elementary fact:
@@
$\#\left\{i=0,\ldots,\left\lfloor\frac{1}{3}\delta(\log n)^{1+\frac{\varepsilon}{8}}\right\rfloor:X_{l_{i+1}}\leq\dfrac{X_{l_{i}}}{(\log n)^{\frac{\varepsilon}{6}}}+(\log n)^{1+\frac{\varepsilon}{2}}\right\}\geq\frac{1}{6}\delta(\log n)^{1+\frac{\varepsilon}{8}},$
then $X_{h^{*}-\lfloor\frac{1}{3}\delta n\rfloor}<2(\log n)^{1+\frac{\varepsilon}{2}}$.